hello, i am completly stuck on this equation i managed i think to work out a part of it but not sure where to go now any help would be greatly appreiciated the equation is 5* 4th sqrt of a -4=3* 4thsqrt of a +2( it didn't have brackets it just had the 5 then the 4th sqrt a-4) i got as far as
5(a^1/4)-4=3a^1/4)+2
then i got stuck

Answer by thederp(1)

5(a^1/4) - 4 = 3(a^1/4) + 2
then it will become:
5(a^1/4) = 3(a^1/4) + 6
5(a^1/4) - 3(a^1/4) = 6
2(a^1/4) = 6
It's just combining like terms. Then, divide everything by 2.
a^1/4 = 3
Raising both sides to the power of 4, you'll get:
a = 81
hope I helped. ^^